How do I express length in feet using mixed fractions, and improper fractions, and a decimal. " a construction worker is using 2 pieces of pipe to form an L-joint for a section of piping that connects a water source to a showerhead. Pipe 1 is 6'4" and Pipe 2 is 8'1". Use this information to answer the following questions.
Express the length of each pipe in feet using mixed fractions.
Express the length of each pipe in feet using improper fractions. Using improper fractions find the length of the L-joint by adding together the lengths of the pipes.

Answer by rfer(16322)

A)
6 1/3 ft
8 1/12 ft
B)
19/3 ft
97/12 ft
C)
19/3+97/12=
76/12+97/12=173/12 ft
reduced
14 5/12 ft total